OpenStudy (anonymous):

i dont really get this. solve each system of equations algebraically y=x+4 y= -x

OpenStudy (anonymous):

if \(y=-x\) and also \(y=x+4\) then since \(y=y\) you know \(-x=x+4\)

OpenStudy (anonymous):

\[-x=x+4\] \[-2x=4\] \[x=-2\]